Circular No. 6467

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

γ-RAY BURST L. Piro, E. Costa, M. Feroci, and M. Cinti, Istituto Astrofisica Spaziale,

CNR, Frascati; F. Frontera, D. Dal Fiume, M. Orlandini, and G. Pizzichini, Istituto Tecnologie e Studio delle Radiazioni Extraterrestri, CNR, Bologna; J. Heise and R. Jager, Space Research Organization of the Netherlands (SRON), Utrecht; and H. Muller, Beppo-SAX Scientific Data Center, Rome, communicate: “On July 20.48395 UT, a γ-ray burst was simultaneously detected in the γ-ray-burst monitor (GRBM, 60–600 keV) and one of the wide-field-camera x-ray detectors (WFC, 2–30 keV) aboard Beppo-SAX. The burst had a peak flux of ∼ 1200 counts/s (2 photons cm^{−2} s^{−1}) in the GRBM, and 90 counts/s (∼ 10^{−8} erg cm^{−2} s^{−1}) in the WFC, with a fluence ratio F (2–30 keV)/F (60–600 keV) of ∼ 0.02. The WFC provided an image of the field with a resolution of 5′ (FWHM). We conservatively estimate a maximum error-box radius of 10′, corresponding to the maximum systematic deviation currently found in position reconstruction. The burst position is α = 17^{h}29^{m}36^{s}, δ = +49^{o}02′.2 (equinox 2000.0). It will be possible to refine the position to a 2′.5 error-box radius after the systematic residuals are calibrated. This position lies within 0^{o}.25 of the center of the 2^{o} error box of a simultaneous BATSE event (trigger #5545, BATSE team, private comm.). A search from all catalogues available via several on-line services for objects within the Beppo-SAX error box yielded no results. The Palomar digitized plates show several sources in the error box, but no one of these is characterized by peculiar colors.”

CAL 83 P. Kahabka, University of Amsterdam; F. Haberl, Max-Planck-Institut

für Extraterrestrische Physik, Garching; and A. N. Parmar, Astrophysics Division, ESTEC, European Space Agency, report: “The supersoft x-ray source CAL 83 has been detected again at a countrate of 0.21 ± 0.04 count/s during a ROSAT/HRI target-of-opportunity observation on Aug. 6.17 UT. The duration of the x-ray off-state on Apr. 28 (IAUC 6432) is limited to < 100 days. This constrains the mass of the white dwarf to ≥ 1.1 M_{⊙} if the off/on behavior is due to an expanding/contracting white-dwarf envelope without mass ejection (cf. Kahabka 1996, in Supersoft X-Ray Sources, ed. J. Greiner, LNP 472, Springer, p. 215).”

1996 September 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green